OpenStudy (anonymous):

cosθ=2/3 and sinθ>0,find cotθ

OpenStudy (anonymous):

Since both sine and cosine are positive, it must be in the first quadrant.

OpenStudy (anonymous):

thanks I agree but what would cotθ be?

OpenStudy (anonymous):

You need to find \(\sin(\theta)\) first.

OpenStudy (anonymous):

Use \[ \cos^2\theta + \sin^2\theta = 1 \]

OpenStudy (anonymous):

cos^ θ+sin^2 θ=1 is the formula but I'm lost.

OpenStudy (anonymous):

Well you know \(\cos \theta = 2/3\) so plug that in.

OpenStudy (anonymous):

\[ \frac 49+\sin^2\theta = 1 \]

OpenStudy (anonymous):

Find \(\sin \theta\) and remember that \(\cot\theta = \cos\theta/\sin\theta\)